Solve and check the following problem using the substution method.
x=1/5Y (this is one over 5)
3x+2y=26

Answer by Alan3354(69443)

Solve and check the following problem using the substution method.
x=1/5Y (this is one over 5) --> x = y/5
3x+2y=26
---------
Sub for x in the 2nd eqn
3*(y/5) + 2y = 26
3y + 10y = 130
13y = 130
y = 10
-----------
x = y/5
x = 2